A 10g 15% gold alloy is mixed with a 20g 5% gold alloy. What is the ercent concentration of the resulting alloy?
Answer by Alan3354(69443)

A 10g 15% gold alloy is mixed with a 20g 5% gold alloy. What is the ercent concentration of the resulting alloy?
-----------
10 gm of 15% = 1.5 gm of gold
20 gm of 5% = 1 gm of gold
2.5 gm of gold in 30 gms total = 2.5/30
= 1/12
=~ 8.33% gold
===========
Or,
10*15 + 20*5 = x*30
30x = 250
x = 8.33%